# Clustering of timed sequences – Application to the analysis of care pathways

### v2

@@ Abstract
keywords:
metric, temporal data, sequential data, medical informatics
journal: Knowledge and Data Engineering
@@ 1 Introduction
Timed sequences, composed of dated events, are types of data frequently encountered in various application domains such as computer log analysis, life course analysis in social science (Gabadinho et al., 2011), tourism trajectories (Bernard et al., 2018) or healthcare pathway analysis (Rama et al., 2019). A specificity of this type of data is that they have no natural vector representation. In particular, they are formed by two very different dimensions: symbolic and temporal; and, on the other hand, their size generally varies from one example to another (e.g., in a healthcare pathway, some patients have more care events than others). This difficulty in modeling data in a vector space makes it challenging to adapt standard data analysis methods (classification, clustering, or forecasting).
@@
The adaptation of data analysis methods for timed sequences raises the question of formalizing a topology of the objects to be analyzed. Most of the data analytic methods make the assumption that objects lie in a metric space. This topology assumption provides intuitive ways to generalize sparse and noisy observations. Formalizing such a space for timed sequences involves defining a metric between timed sequences. This problem is very similar to one encountered in time series analysis (temporal data where values are numerical, not event-based, and generally with a regular and dense temporal sampling). For time series analysis, a large collection of metrics has been proposed and allows them to be analyzed (Holder et al., 2024; Tavenard et al., 2020). The challenges turns out to identify a metric that has the properties appropriate for the analysis task one wishes to carry out (e.g., the properties of a distance) and whose semantics will give the expected results. Intuitively, a similarity metric (resp. dissimilarity metric) should produce a high similarity value (resp. low dissimilarity value) when two examples are considered similar by domain experts.
@@
For sequential data, few studies have focused on adapting data analysis methods. Gay et al. (2015) and H. Sebia (2024) are based on tensor representation of sequential data, while Yu et al. (2020) or Nguyen et al. (2018) use embedding representations. Nonetheless, only a few methods handle event logs with timestamped events dates (Rama et al., 2019; Gabadinho et al., 2011). The TraMineR (Gabadinho et al., 2011) software has been frequently used for epidemiological studies, for instance by Le Meur et al. (2019). It enables the user to cluster or classify event-based temporal data with different metrics. Nonetheless, this method represents an event-based data as a state-sequence, i.e., a longitudinal sequence which associate at each instant a unique state. One advantage of this tool is to propose an end-to-end analysis of the data. Another end-to-end method for timed sequences is AliClu (Rama et al., 2019) which combine the definition of a time warping metric with a hierarchical clustering algorithm. Contrary to TraMineR, it does not allow the change of the metric. Beside these end-to-end methods, some works were dedicated to metrics between temporal sequences (Garreau et al., 2014; Schneider et al., 2022; Compagnon et al., 2021) or averaging (Chen et al., 2017; Besnard and Guyet, 2023; Tatti and Vreeken, 2012). Despite these proposals, metrics whose semantics can better take into account the temporal dimension are still lacking. In particular, the delays between the occurrence of events is not handle. Most current solutions generally only take into account the sequentiality in the data.
@@
In this work, we extend the drop-DTW metric (Dvornik et al., 2021) which was proposed to evaluate distances between sequences by taking into account both sequential aspects, by adapting Dynamic Time Warping (Sakoe and Chiba, 1978), and by adding support for meaningless events, which can be dropped. This metric has the semantics that seem appropriate for our application case – i.e., the clustering of care pathways, but as this metric is not a distance, it does not allow for the construction of an average sequence nor to be used in classical clustering methods such as the K-Means algorithm or hierarchical clustering.
@@
The main contribution of this article is the proposal of an algorithm to compute a mean timed sequence based on drop-DTW. This algorithm is inspired by the DBA method (Petitjean et al., 2011) for time series. The construction of an average sequence meets both the needs of clustering algorithms (hierarchical clustering or K-means) and the needs of interpretation of sequence clusters (definition of a representative).
@@
This method is applied on a Electronic Health Records (EHR) dataset in order to discover typical care-pathways. In the context of the OPTISOINS project [^1] https://www.bernoulli-lab.fr/project/optisoins/, we were particularly interested in patients who had undergone pulmonary resection surgery for oncological reasons. Care trajectories were reconstructed from data collected in the EHR data warehouse of the Greater Paris Hospital (AP-HP). The care pathway is a timed sequence representing the key stages of managing the patient surgery from diagnosis to post-surgery treatments.
@@
The reminder of this article is organized as follow. We first introduce the notations related to timed sequences. Then, Section 3 introduces the adaptation of drop-DTW to timed sequences. Section 4 presents the adaptation of the DBA algorithm for averaging timed sequences based on the drop-DTW. This averaging method is required in classical clustering algorithms that are briefly recalled in Section 5. Finally, Sections 6 and 7 present experiments and results on synthetic data and our use case (OPTISOINS project) respectively.
@@ 2 Times Sequences and Probabilistic Timed Sequences
In this section, we introduce the notations for timed sequences and probabilistic timed sequences. The probabilistic version of timed sequences enables to adapt the metrics defined for time series to timed sequences (see next section).
@@ 3.1 DTW and drop-DTW
Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) computes the optimal alignment between two series subject to certain constraints. Here, we adapt these definitions to the case of timed sequences. Let $X=\langle(x_{1},t^{x}_{1}),\dots,(x_{N},t^{x}_{N})\rangle$ and $Z=\langle(z_{1},t^{z}_{1})\dots,(z_{K},t^{z}_{K})\rangle$ be two input sequences, where $N,K$ are their respective lengths. An alignment between two sequences is defined as a binary matrix, $\pi\in\{0,1\}^{K\times N}$ where $\pi_{i,j}=1$ if $z_{i}$ is matched to $x_{j}$, and 00 otherwise. Matching an element $z_{i}$ to an element $x_{j}$ incurs a cost $C_{i,j}$ between the $i$-th and $j$-th elements of the sequences. The cost values are user-defined. DTW finds the alignment $\pi^{*}$ between sequences $Z$ and $X$ that minimizes the global cost:
@@
drop-DTW (Dvornik et al., 2021) extends the set of feasible alignments, $\mathcal{P}$, to those adhering only to the monotonicity constraint. Consequently, unlike DTW, events can be dropped from the alignment process. drop-DTW finds the optimal temporal alignment while allowing for the removal of certain elements. To account for unmatched elements in the alignment cost calculation, Dvornik et al. (2021) introduce a new drop cost: $\delta\in\mathbb{R}^{+}$. [^2] In the original approach of Dvornik et al., the cost value depends on the event type. We use a constant value for the sake of simplification. The optimal matching can then be defined as follows:
@@ 3.2 Adaptation of drop-DTW for timed sequences
The proposed adaptation of drop-DTW focuses on proposing a metric between events and adding temporal constraints to the drop-DTW algorithm. The Dynamic Time Warping algorithm requires defining a cost between two events to compare sequences. To take into account both time and events, we decided to assign weights to a time distance cost $c_{t}$, and an event difference cost $c_{e}$. Having introduced the probabilistic representation of timed sequences, we define the distance between two events $e_{a}=(\bm{d}_{a}=[d^{1}_{a},\dots,d^{n}_{a}],t_{a})$ and $e_{b}=(\bm{d}_{b}=[d^{1}_{b},\dots,d^{n}_{b}],t_{b})$ by:
@@
where $p_{t},p_{e}\in\mathbb{R}_{+}$ are weights of the two terms: a distance between probabilistic events and a time gap. This metric implies that time events are in an Euclidean space of dimension $n+1$. This is of particular interest in the next section, to define average events.
@@ 4 Construction of an Average Timed Sequence
Computing an average timed sequence consistent with the notion of similarity is essential for clustering algorithms. In the following, we propose a method inspired by the DBA algorithm (Petitjean and Gançarski, 2012) to create average sequences in terms of drop-DTW, taking into account the possibility of removal. DBA addresses the problem of generating an average time series in terms of DTW – which does not hold distance properties. In this case, the generalized average is defined as a time series that minimizes the global inertia (see Equation 1).
@@ 4.1 Average Timed Sequences
Let $\mathcal{S}=\{s_{1},\dots,s_{m}\}$ be a set of timed sequences, Algorithm 2 details the steps for generating an average timed sequence inspired by the principle of vertical averages of DBA (Petitjean et al., 2011). In this algorithm, $\mathcal{C}^{(k)}$ denotes the current average sequence at the $k$-th iteration. At line 1, the current average time sequence, $\mathcal{C}^{(0)}$ is initialized with a timed sequence chosen from $\mathcal{S}$. Then, this current average is processed into the outer loop which repeats $maxit$ times the alignment and update steps. The alignment step (line 6-11) calculates the alignment by drop-DTW between $\mathcal{C}^{(0)}$ and each of the other sequences. The alignment to one sequence is obtained line 8 from the distance matrix computed the drop-DTW. We refer the reader to the original DTW algorithm to get the details on how to compute an alignment from a distance matrix. Lines 9 to 11 collect, on the one hand, the list of paired events in the sequences of $\mathcal{S}$ per event id, $e$, in $\mathcal{E}_{e}$ and, on the other hand, their dates per event id, $e$, in $\mathcal{T}_{e}$. These lists are then averaged lines 12-13. This is the update step. This vertical average of events is constructed as the histogram of event distributions. Lastly, line 14 creates a new average sequence which can be refined in a subsequent iteration until convergence. In our implementation, we assume convergence after a fixed number of iterations but a convergence criteria could be implemented to prevent useless iterations.
Similar to DBA, events from probabilistic sequences are represented in a vector space. Thus, the new average sequence is always closer (according to drop-DTW) to the sequences it averages, and the algorithm converges well (see Section 4.2). Additionally, this algorithm preserves the drop-DTW’s ability to allow temporal shifts and handle missing or aberrant events.
The choice of the timed sequence in line 1 is crucial for the result. We choose to randomly select a sequence of the longest length. We made this choice because our algorithm decreases the length of the current average timed sequence, but it can not increase it. Note that there are some smarter initialization strategies for k-means like algorithms that could improve the convergence, for instance K-means++ (Arthur and Vassilvitskii, 2007), or recent provably good seedings methods (Bachem et al., 2016).
@@ 8       $k\leftarrow k+1$;
Figure 2 illustrates one iteration of average calculation for the three sequences of Example 1. It assumes that we already have a current average sequence $s_{r}$ (a surgery, then a consultation and two radiotherapy sessions). The figure depicts the alignments by drop-DTW of $s_{r}$ with each sequence of Example 1. The gray lines represents the alignments (line 8). $s^{\prime}_{r}$ is the updated average sequence obtained by the lines 11-12 of Algorithm 2. As the $C$ event has been paired with both $C$ and $R$ events, the second probabilistic event of $s^{\prime}_{r}$ combines the two types of events (proportionally). In addition, the first $R$ event of $s_{r}$ is removed from $s^{\prime}_{r}$ because it is paired with no event. The other events are purely paired with similar event types, then the probabilistic event is a pure $S$ or $R$ in $s^{\prime}_{r}$ but their time positions have been adjusted with respect to the positions of the new aligned events.
@@ 4.2 Convergence of the Algorithm
The algorithm 2 refines iteratively an average sequence for a set of sequences, $\left\{s_{i}\right\}_{i\in[m]}$. At step $k$, the inertia of the average sequence is defined as:
@@
Let us not have a look at the sum over $i$ of the $q_{i}^{(k+1)}$. This sum can be rewritten as follow:
@@ 4.3 Parameters
The ratio $\frac{p_{t}}{p_{e}}$ can be inferred from the delay $\tau$ by the following reasoning. For events located within the time interval $[t-\tau,t+\tau]$, the importance of the event type should be greater than the importance of time. We must then choose $p_{e}$ to be at least $\tau^{2}$ times greater than $p_{t}$. Without lost of generality, the value of $p_{t}$ can be set at $1$ by default. The drop-cost is also derived from $\tau$, i.e., the delay from which there is no longer any interest in associating two events. We thus propose to define the drop-cost as:
@@ 5 Clustering of Timed Sequences
The proposed metric between timed sequences and the averaging techniques are used to adapt two classical clustering techniques: a hierarchical ascendant clustering (HAC) and a K-means algorithm (Hartigan, 1975). These two algorithms require to predefine the expected number of clusters, $K$ [^3] Model selection based on the inertia of the clustering may be applied to select the optimal number of clusters.
@@
The $K$-means also proceeds iteratively to refine $K$ clusters. K-means algorithm has two main steps: first the sequences of each cluster are averaged (Algorithm 2). The computed averages are the representatives of their clusters. Then, new clusters are made by assigning optimally each sequence to one of the $K$ representatives.
@@ 6.1 Effects of the $\frac{p_{t}}{p_{e}}$ Ratio
In the case of the first configuration, we observe that all clusters generated by HierAsTiSeq were correctly identified (Kappa score of $1$). However, in the case of the second configuration, we observe a degradation of results with a Kappa score of $0.62$. This degradation was expected because the maximum time distance observed between events are around 20. Consequently, the notion of time is practically never taken into account by the metric.

### v3

# Thomas Guyet, Pierre Pinson and Enoal Gesny†,‡

@@ 1 Introduction
Timed sequences, composed of dated events, are types of data frequently encountered in various application domains such as computer log analysis, life course analysis in social science [8], tourism trajectories [3] or healthcare pathway analysis [18]. A specificity of this type of data is that they have no natural vector representation. In particular, they are formed by two very different dimensions: symbolic and temporal; and, on the other hand, their size generally varies from one example to another (e.g., in a healthcare pathway, some patients have more care events than others). This difficulty in modeling data in a vector space makes it challenging to adapt standard data analysis methods (classification, clustering, or forecasting).
@@
The adaptation of data analysis methods for timed sequences raises the question of formalizing a topology of the objects to be analyzed. Most data analysis methods make the assumption that objects lie in a metric space. This topology assumption provides intuitive ways to generalize sparse and noisy observations. Formalizing such a space for timed sequences involves defining a metric between timed sequences. This problem is very similar to the one encountered in time series analysis (temporal data where the values are numerical, not event-based and generally with a regular and dense temporal sampling). For time series analysis, a large collection of metrics has been proposed and allows them to be analyzed [13, 23]. The challenges reveals to identify a metric that has the appropriate properties for the analysis task one wishes to carry out (e.g., the properties of a distance) and whose semantics will give the expected results. Intuitively, a similarity metric (resp. dissimilarity metric) should produce a high similarity value (resp. low dissimilarity value) when two examples are considered similar by domain experts.
For sequential data, few studies have focused on adapting data analysis methods. [10] and [11] are based on tensor representation of sequential data, while [24] or [15] use embedding representations. Nonetheless, only a few methods handle event logs with timestamped events dates [18, 8].
@@
The TraMineR [8] software has been frequently used for epidemiological studies, for instance by [14]. It enables the user to cluster or classify event-based temporal data with different metrics. Nonetheless, this method represents an event-based data as a state-sequence, i.e., a longitudinal sequence which associate at each instant a unique state. One advantage of this tool is to propose an end-to-end analysis of the data. Another end-to-end method for timed sequences is AliClu [18] which combines the definition of a time warping metric with a hierarchical clustering algorithm. Contrary to TraMineR, it does not allow a change of metric.
In addition to these end-to-end methods, some work was dedicated to metrics between temporal sequences [9, 21, 6] or averaging [5, 4, 22]. Despite these proposals, metrics whose semantics can better take into account the temporal dimension are still lacking. In particular, the delays between the occurrence of events are not handled. Most current solutions generally only take into account the sequentiality in the data.
@@
In this work, we extend the drop-DTW metric [7] which was proposed to evaluate distances between sequences taking into account both sequential aspects, by adapting Dynamic Time Warping [20], and by adding support for meaningless events that can be dropped. This metric has the semantics that seem appropriate for our application case – i.e., the clustering of care pathways, but as this metric is not a distance, it does not allow for the construction of an average sequence nor to be used in classical clustering methods such as the K-Means algorithm or hierarchical clustering.
@@
The main contribution of this article is the proposal of an algorithm to compute a mean timed sequence based on drop-DTW. This algorithm is inspired by the DBA method [17] for time series. The construction of an average sequence meets both the needs of clustering algorithms (hierarchical clustering or K-means) and the needs of interpretation of sequence clusters (definition of a representative).
@@
This method is applied on a Electronic Health Records (EHR) dataset in order to discover typical care-pathways. In the context of the OPTISOINS project [^2] https://www.bernoulli-lab.fr/project/optisoins/, we were particularly interested in patients who had undergone pulmonary resection surgery for oncological reasons. Care trajectories were reconstructed from data collected in the EHR data warehouse of the Greater Paris Hospital (AP-HP). The care pathway is a timed sequence representing the key stages of managing patient surgery from diagnosis to post-surgery treatments.
@@
The reminder of this article is organized as follows. We first introduce notations related to timed sequences. Then, Section 3 introduces the adaptation of drop-DTW to timed sequences. Section 4 presents the adaptation of the DBA algorithm for averaging timed sequences based on the drop-DTW. This averaging method is required in classical clustering algorithms that are briefly recalled in Section 5. Finally, Sections 6 and 7 present experiments and results on synthetic data and our use case (OPTISOINS project) respectively.
@@ 2 Times Sequences and Probabilistic Timed Sequences
In this section, we introduce notations for timed sequences and probabilistic timed sequences. The probabilistic version of timed sequences enables to adapt the metrics defined for time series to timed sequences (see next section).
@@ 3.1 DTW and drop-DTW
Dynamic Time Warping (DTW) [20] computes the optimal alignment between two series subject to certain constraints. Here, we adapt these definitions to the case of timed sequences. Let $X=\langle(x_{1},t^{x}_{1}),\dots,(x_{N},t^{x}_{N})\rangle$ and $Z=\langle(z_{1},t^{z}_{1})\dots,(z_{K},t^{z}_{K})\rangle$ be two input sequences, where $N,K$ are their respective lengths. An alignment between two sequences is defined as a binary matrix, $\pi\in\{0,1\}^{K\times N}$ where $\pi_{i,j}=1$ if $z_{i}$ is matched to $x_{j}$, and 00 otherwise. Matching an element $z_{i}$ to an element $x_{j}$ incurs a cost $C_{i,j}$ between the $i$-th and $j$-th elements of the sequences. The cost values are user-defined. DTW finds the alignment $\pi^{*}$ between sequences $Z$ and $X$ that minimizes the global cost:
@@
drop-DTW [7] extends the set of feasible alignments, $\mathcal{P}$, to those adhering only to the monotonicity constraint. Consequently, unlike DTW, events can be dropped from the alignment process. drop-DTW finds the optimal temporal alignment while allowing for the removal of certain elements. To account for unmatched elements in the alignment cost calculation, [7] introduce a new drop cost: $\delta\in\mathbb{R}^{+}$. [^3] In the original approach of Dvornik et al., the cost value depends on the event type. We use a constant value for the sake of simplification. The optimal matching can then be defined as follows:
@@ 3.2 Adaptation of drop-DTW for timed sequences
The proposed adaptation of drop-DTW focuses on proposing a metric between events and adding temporal constraints to the drop-DTW algorithm. The Dynamic Time Warping algorithm requires defining a cost between two events to compare sequences. To take into account both time and events, we decided to assign weights to a time distance cost $c_{t}$, and an event difference cost $c_{e}$. Having introduced the probabilistic representation of timed sequences, the distance between two events $e_{a}=(\bm{d}_{a}=[d^{1}_{a},\dots,d^{n}_{a}],t_{a})$ and $e_{b}=(\bm{d}_{b}=[d^{1}_{b},\dots,d^{n}_{b}],t_{b})$ is defined by:
@@
where $p_{t},p_{e}\in\mathbb{R}_{+}$ are weights of two terms: a distance between probabilistic events and a time gap. This metric implies that time events are in an Euclidean space of dimension $n+1$. This is of particular interest in the next section, to define average events.
@@ 4 Construction of an Average Timed Sequence
Computing an average timed sequence consistent with the notion of similarity is essential for clustering algorithms. In the following, we propose a method inspired by the DBA algorithm [16] to create average sequences in terms of drop-DTW, taking into account the possibility of removal. DBA addresses the problem of generating an average time series in terms of DTW – which does not hold distance properties. In this case, the generalized average is defined as a time series that minimizes the global inertia (see Equation 1).
@@ 4.1 Average Timed Sequences
Let $\mathcal{S}=\{s_{1},\dots,s_{m}\}$ be a set of timed sequences, Algorithm 2 details the steps for generating an average timed sequence inspired by the principle of vertical averages of DBA [17]. In this algorithm, $\mathcal{C}^{(k)}$ denotes the current average sequence at the $k$-th iteration. At line 1, the current average time sequence, $\mathcal{C}^{(0)}$ is initialized with a timed sequence chosen from $\mathcal{S}$. Then, this current average is processed into the outer loop which repeats $maxit$ times the alignment and update steps. The alignment step (lines 6-11) calculates the alignment by drop-DTW between $\mathcal{C}^{(0)}$ and each of the other sequences. The alignment to one sequence is obtained line 8 from the distance matrix computed the drop-DTW. We refer the reader to the original DTW algorithm to get the details on how to compute an alignment from a distance matrix. Lines 9 to 11 collect, on the one hand, the list of paired events in the sequences of $\mathcal{S}$ per event id, $e$, in $\mathcal{E}_{e}$ and, on the other hand, their dates per event id, $e$, in $\mathcal{T}_{e}$. These lists are then averaged lines 12-13. This is the update step. This vertical average of events is constructed as the histogram of event distributions. Lastly, line 14 creates a new average sequence which can be refined in a subsequent iteration until convergence. In our implementation, we assume convergence after a fixed number of iterations but a convergence criteria could be implemented to prevent useless iterations.
Similarly to DBA, events from probabilistic sequences are represented in a vector space. Thus, the new average sequence is always closer (according to drop-DTW) to the sequences it averages, and the algorithm converges well (see Section 4.2). Additionally, this algorithm preserves the drop-DTW’s ability to allow temporal shifts and handle missing or aberrant events.
The choice of the timed sequence in line 1 is crucial for the result. We choose to randomly select a sequence of the longest length. We made this choice because our algorithm decreases the length of the current average timed sequence, but it can not increase it. Note that there are some smarter initialization strategies for k-means like algorithms that could improve the convergence, for instance K-means++ [1], or recent provably good seeding methods [2].
@@ 8       $k\leftarrow k+1$;
Figure 2 illustrates one iteration of average calculation for the three sequences in Example 1. It assumes that we already have a current average sequence $s_{r}$ (a surgery, then a consultation and two radiotherapy sessions). The figure depicts the alignments by drop-DTW of $s_{r}$ with each sequence of Example 1. The gray lines represents the alignments (line 8). $s^{\prime}_{r}$ is the updated average sequence obtained by the lines 11-12 of Algorithm 2. As the $C$ event has been paired with both $C$ and $R$ events, the second probabilistic event of $s^{\prime}_{r}$ combines the two types of events (proportionally). In addition, the first $R$ event of $s_{r}$ is removed from $s^{\prime}_{r}$ because it is paired with no event. The other events are purely paired with similar event types, then the probabilistic event is a pure $S$ or $R$ in $s^{\prime}_{r}$ but their time positions have been adjusted with respect to the positions of the new aligned events.
@@ 4.2 Convergence of the Algorithm
The algorithm 2 iteratively refines an average sequence for a set of sequences, $\left\{s_{i}\right\}_{i\in[m]}$. At step $k$, the inertia of the average sequence is defined as:
@@
Let us not have a look at the sum over $i$ of the $q_{i}^{(k+1)}$. This sum can be rewritten as follows:
@@ 4.3 Parameters
The ratio $\frac{p_{t}}{p_{e}}$ can be inferred from the delay $\tau$ by the following reasoning. For events located within the time interval $[t-\tau,t+\tau]$, the importance of the event type should be greater than the importance of time. We must then choose $p_{e}$ to be at least $\tau^{2}$ times greater than $p_{t}$. Without lost of generality, the value of $p_{t}$ can be set at $1$ by default. The drop-cost is also derived from $\tau$, i.e., the delay from which there is no longer any interest in associating two events. Thus, we propose to define the drop-cost as:
@@ 5 Clustering of Timed Sequences
The proposed metric between timed sequences and the averaging technique are used to adapt two classical clustering techniques: a hierarchical ascendant clustering (HAC) and a K-means algorithm [12]. These two algorithms require to predefine the expected number of clusters, $K$ [^4] Model selection based on the inertia of the clustering may be applied to select the optimal number of clusters.
@@
The $K$-means algorithm also proceeds to refine iteratively $K$ clusters. K-means algorithm has two main steps: first the sequences of each cluster are averaged (Algorithm 2). The computed averages are the representatives of their clusters. Then, new clusters are made by assigning optimally each sequence to one of the $K$ representatives.
@@ 6.1 Effects of the $\frac{p_{t}}{p_{e}}$ Ratio
In the case of the first configuration, we observe that all clusters generated by HierAsTiSeq were correctly identified (Kappa score of $1$). However, in the case of the second configuration, we observe a degradation of results with a Kappa score of $0.62$. This degradation was expected because the maximum time distance observed between events is around 20. Consequently, the notion of time is practically never taken into account by the metric.